# Cosmic Ray Muons in Laboratories Deep Underground

### v1

@@ VI Underground Spectra and Mean Energies
Figure 9: Underground muon energy spectra for labs under flat overburdens and mountains, calculated by mute with Eq. 15. The energy spectrum given in Ref. Tang et al. (2006) for Kamioka from a music/musun simulation is shown by the thick grey curve.
@@
The mean underground muon energy is also relevant to underground detectors, particularly in stopping vs through-going muons studies. Historically, it has been calculated by means of a parameterized energy spectrum,
@@
where $A$ is a normalization constant with respect to the differential muon intensity at a given depth $h$ Gaisser et al. (2016); Workman et al. (2022); Mei and Hime (2006). Values for the parameters $\epsilon_{\mu}$, $b$, and $\gamma_{\mu}$ for standard rock are listed in Ref. Mei and Hime (2006) for the parameterizations from both Lipari, et al. Lipari and Stanev (1991) ($\epsilon_{\mu}=618~{}\mathrm{GeV},b=0.383~{}\mathrm{km.w.e.}^{-1},\gamma_{\mu}=3.7$), and Groom, et al. Groom et al. (2001); Battistoni et al. (1999) ($\epsilon_{\mu}=693~{}\mathrm{GeV},b=0.4~{}\mathrm{km.w.e.}^{-1},\gamma_{\mu}=3.77$).
@@ Acknowledgements.
We acknowledge the help of Marco Selvi and the LVD collaboration, who provided us with the topographic map of the Gran Sasso mountain and the LVD data for the muon angular distribution. We acknowledge, as well, the help of Michel Zampaolo and Luigi Mosca for the data of the Fréjus detector, and Holger Kluck and Wolfgang Rhode for a map of the Fréjus mountain. We also acknowledge Eunju Jeon and Hyun Su Lee, who provided us with a topographic map of the Yangyang Mountain and discussed with us muon bundles in the COSINE-100 detector, Itaru Shimizu and Shigetaka Moriyama for the maps of the Kamioka Mountain, and Shaomin Chen for the map of the Jinping Mountain. We want to thank Joseph Formaggio, Christopher Kyba, and Chris Waltham for their helpful discussions about the SNO detector and measurement. In addition, we acknowledge Jaret Heise and Kyle Jankord for providing the geochemical information for SURF rock formations and a three-dimensional SURF geology model, as well as the help of Joshua Albert for information on the overburden of WIPP. This research was enabled in part by support provided by the Digital Research Alliance of Canada and the Academia Sinica Grid-Computing Center. W. Woodley and M.-C. Piro acknowledge the support from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute.
@@ Appendix B Values Used to Model Laboratory Rock Types
As described in Sections II and III, rock compositions and Sternheimer parameters for different media are used in our proposal simulations to calculate energy losses in mute v3. Table 5 and Table 6 report the minor component chemical compositions and Sternheimer parameters, respectively, which are used to obtain the main results in our work. The $Z/A\right>$ and $Z^{2}/A\right>$ values are defined as:
@@
Table 5: Percent weights (%) of minor component rock compositions for laboratories considered in this study, where available. Rock compositions were unavailable for WIPP, Boulby, and SUPL. In some cases, percent weights might not sum to 100%, as some minor components that do not contribute significantly to the make-up of the rock have been excluded from reporting.
@@
Kamioka [^9] Multiple rock samples from the Kamioka mountain are listed in Ref. Mizukoshi et al. (2018). Here we have used the average of the two JR-1 and JA-3 igneous samples, which are widely distributed around the Kamioka area, as given in Ref. Yoon et al. (2021). 0 0 39.00 0.01 0.60 6.00 17.00 0 28.00 7.60 Yoon et al. (2021); Mizukoshi et al. (2018)
@@
Table 6: Sternheimer parameters for media used in this work. $I$ is the ionisation energy of the medium, and the $\delta_{0}$ parameter from Eq. 8 is 0 for all media. Values are taken from Koehne et al. (2013).
@@
The definitions of freshwater and seawater are very similar in terms of their Sternheimer parameters. However, their mass densities — given in Section II — and their chemical compositions are different Koehne et al. (2013). These differences lead to total fluxes calculated with sea water being lower than those calculated with fresh water by about 5%, constant across all depths.

### v2

@@ VI Underground Spectra and Mean Energies
Figure 9: Underground muon energy spectra for labs under flat overburdens and mountains, calculated by mute with Eq. 15. The energy spectrum given in Ref. Tang et al. (2006) for Kamioka from a music/musun simulation is shown by the thick gray curve.
@@
The mean underground muon energy is also relevant to underground detectors, particularly in stopping vs through-going muons studies. Historically, it has been calculated by means of a parametrized energy spectrum,
@@
where $A$ is a normalization constant with respect to the differential muon intensity at a given depth $h$ Gaisser et al. (2016); Workman et al. (2022); Mei and Hime (2006). Values for the parameters $\epsilon_{\mu}$, $b$, and $\gamma_{\mu}$ for standard rock are listed in Ref. Mei and Hime (2006) for the parametrizations from both Lipari, et al. Lipari and Stanev (1991) ($\epsilon_{\mu}=618~{}\mathrm{GeV},b=0.383~{}\mathrm{km.w.e.}^{-1},\gamma_{\mu}=3.7$), and Groom, et al. Groom et al. (2001); Battistoni et al. (1999) ($\epsilon_{\mu}=693~{}\mathrm{GeV},b=0.4~{}\mathrm{km.w.e.}^{-1},\gamma_{\mu}=3.77$).
@@ Acknowledgements.
We acknowledge the help of Marco Selvi and the LVD collaboration, who provided us with the topographic map of the Gran Sasso mountain and the LVD data for the muon angular distribution. We acknowledge, as well, the help of Michel Zampaolo and Luigi Mosca for the data of the Fréjus detector, and Holger Kluck and Wolfgang Rhode for a map of the Fréjus mountain. We also acknowledge Eunju Jeon and Hyun Su Lee, who provided us with a topographic map of the Yangyang Mountain and discussed with us muon bundles in the COSINE-100 detector, Itaru Shimizu and Shigetaka Moriyama for the maps of the Kamioka Mountain, and Shaomin Chen for the map of the Jinping Mountain. We want to thank Joseph Formaggio, Christopher Kyba, and Chris Waltham for their helpful discussions about the SNO detector and measurement. In addition, we acknowledge Jaret Heise and Kyle Jankord for providing the geochemical information for SURF rock formations and a three-dimensional SURF geology model, as well as the help of Joshua Albert for information on the overburden of WIPP. This research was enabled in part by support provided by the Digital Research Alliance of Canada and the Academia Sinica Grid-Computing Center (Grant No. AS-CFII-112-103). W. Woodley and M.-C. Piro acknowledge the support from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute.
@@ Appendix B VALUES USED TO MODEL LABORATORY ROCK TYPES
As described in Secs. II and III, rock compositions and Sternheimer parameters for different media are used in our proposal simulations to calculate energy losses in mute v3. Tables 5 and 6 report the minor component chemical compositions and Sternheimer parameters, respectively, which are used to obtain the main results in our work. The $Z/A\right>$ and $Z^{2}/A\right>$ values are defined as:
@@
Table 5: Percent weights (%) of minor component rock compositions for laboratories considered in this study, where available. Rock compositions were unavailable for WIPP, Boulby, and SUPL. In some cases, percent weights might not sum to 100%, as some minor components that do not contribute significantly to the makeup of the rock have been excluded from reporting.
@@
Kamioka [^9] Multiple rock samples from Ikenoyama Mountain above the Kamioka lab are listed in Ref. Mizukoshi et al. (2018). Here we have used the average of the two JR-1 and JA-3 igneous samples, which are widely distributed around the Kamioka area, as given in Ref. Yoon et al. (2021). 0 0 39.00 0.01 0.60 6.00 17.00 0 28.00 7.60 Yoon et al. (2021); Mizukoshi et al. (2018)
@@
Table 6: Sternheimer parameters for media used in this work. $I$ is the ionization energy of the medium, and the $\delta_{0}$ parameter from Eq. 8 is 0 for all media. Values are taken from Koehne et al. (2013).
@@
The definitions of fresh water and sea water are very similar in terms of their Sternheimer parameters. However, their mass densities—given in Sec. II—and their chemical compositions are different Koehne et al. (2013). These differences lead to total fluxes calculated with sea water being lower than those calculated with fresh water by about 5%, constant across all depths.